Apply the quadratic formula to find the roots of the given function, and then graph the function g(x) = x^2 + x + 12
Answer by rfer(16322)

f(x)=2^2+2+12=18 (2,18)
f(x)=-2^2+-2+12=14 (-2,14)
f(x)=4^2+4+12=32 (4,32)
f(x)=-4^2+-4+12=24 (-4,24)
vertex=(0,12)